# BRcal: An R Package to Boldness-Recalibrate Probability Predictions

### v2

@@ Abstract
When probability predictions are too cautious for decision making, boldness-recalibration enables responsible emboldening while maintaining the probability of calibration required by the user. We introduce BRcal, an R package implementing boldness-recalibration and supporting methodology as recently proposed. The BRcal package provides direct control of the calibration-boldness tradeoff and visualizes how different calibration levels change individual predictions. We describe the implementation details in BRcal related to non-linear optimization of boldness with a non-linear inequality constraint on calibration. Package functionality is demonstrated via a real world case study involving housing foreclosure predictions. The BRcal package is available on the Comprehensive R Archive Network (CRAN) (https://cran.r-project.org/web/packages/BRcal/index.html) and on Github (https://github.com/apguthrie/BRcal).
@@
Acknowledgements The authors are thankful to Leidos for providing funding for this work. The authors would like to thank Matthew Keefe and Andrew McCoy for their insights and roles obtaining case study data for this line of research.
@@ 1 Introduction
Boldness-recalibration is a technique that responsibly emboldens probability predictions of binary events while maintaining a user chosen posterior probability of calibration [1]. When events occur at the rates that were predicted, the probability predictions are well calibrated. Calibration is important so decision makers can understand the long run rates of events. Overly cautious predictions tend to satisfy the literal definition of calibration (e.g. predicting sample proportion $\bar{y}$ for all events). When predictions exhibit greater spread, or “boldness”, they are more actionable. Boldness-recalibration leverages this tradeoff between calibration and boldness by maximizing boldness at a specified level of calibration, subject to classification accuracy. Here, calibration is measured using the posterior model probability of calibration. The posterior model probability of calibration is easily interpretable in the context of an individual forecaster, which is an advantage for a calibration metric. The key feature of boldness-recalibration that sets it apart from most other calibration techniques is that it provides responsible emboldening of predictions rather than just improving calibration. In this paper, we present BRcal: an R package implementing boldness-recalibration and supporting methodology including calibration assessment via Bayesian model selection, LLO-adjustment via MLEs, and visualizations.
@@
The rest of this paper is organized as follows. Section 2 briefly reviews the methodology behind boldness-recalibration and related techniques. Section 3 introduces the BRcal package by describing its core functions and implementation. In Section 4, we demonstrate the capabilities of BRcal by presenting an example related to housing foreclosure predictions. Section 5 provides concluding remarks.
2 Methods
We introduced boldness-recalibration, a forecaster agnostic approach, to improve upon established recalibration techniques in situations where overly cautious predictions hinder the decision making process [1]. However, it is important to assess calibration and boldness before boldness-recalibrating. Rather than rely on calibration assessment techniques such as the Brier Score, Expected Calibration Error (ECE), and Log Loss, we employ a Bayesian model selection approach to measure the probability of calibration. Throughout this paper and in the BRcal package, we focus on probability predictions of binary events.
2.1 Background
For a more complete exposition of methods, see [1]. A brief and self-contained review is found here. Let $\mathbf{x}$ be a vector of probability predictions from a single source (i.e. a single forecaster, or a single machine learning model, etc) and let $\mathbf{y}$ be a vector of binary event outcomes. Following the convention of [1], we adopt the following likelihood function
$\pi(\mathbf{x},\mathbf{y}|\delta,\gamma)=\prod_{i=1}^{n}c(x_{i};\delta,\gamma)^{y_{i}}\left[1-c(x_{i};\delta,\gamma)\right]^{1-y_{i}}.$ (1)
@@ 2.1 Background
and $\delta>0$ and $\gamma\in\mathbb{R}$ are recalibration parameters [23, 24]. Equation (2) maps probability forecasts to the log odds scale, shifts them by log $\delta$, scales by $\gamma$, and remaps back to the probability scale. When we apply (2) to prediction $x_{i}$, we call this “LLO-adjustment”.
To assess calibration, we take an approximate Bayesian model selection-based approach comparing a well calibrated model ($\delta$ = $\gamma$ = 1) denoted by $M_{c}$, to an uncalibrated model ($\delta>0$ and $\gamma\in\mathbb{R}$) denoted by $M_{u}$. The posterior model probability of $M_{c}$ is defined as
$P(M_{c}|\mathbf{y})=\frac{1}{1+BF\frac{P(M_{u})}{P(M_{c})}}.$ (3)
where $BF$ is the Bayes Factor comparing $M_{u}$ to $M_{c}$ defined as
$BF=\frac{P(\mathbf{y}|M_{u})}{P(\mathbf{y}|M_{c})}.$ (4)
@@
$BF\approx exp\left\{-\frac{1}{2}(BIC_{u}-BIC_{c})\right\}.$ (5)
Here, the BIC under the well calibrated model $M_{c}$ is defined as
@@
$BIC_{u}$ $=2\times log(n)-2\times log(\pi(\hat{\delta}_{MLE},\hat{\gamma}_{MLE}|\mathbf{x},\mathbf{y})).$ (7)
The quantity $P(M_{c}|\mathbf{y})$ is the probability $\mathbf{x}$ is calibrated given observed outcomes $\mathbf{y}$ and serves as our measure of calibration.
To recalibrate $\mathbf{x}$ such that calibration is maximized, we use (1) to obtain maximum likelihood estimates (MLEs) for the recalibration parameters, denoted $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$. Then, these MLEs are used to obtain the maximally calibrated predictions $\mathbf{x}_{MLE}$ where $x_{i,MLE}=c(x_{i};\hat{\delta}_{MLE},\hat{\gamma}_{MLE})$. We call this approach MLE-recalibration.
2.2 Boldness-Recalibration
Maximally calibrated predictions are not necessarily bold enough for decision making. For example, the Hockey study in [1] showed that even after MLE-recalibration, the expert forecaster’s predictions could be more bold while maintaining a high probability of calibration via boldness-recalibration. On the other hand, random noise predictions were reigned in close to the base rate of 0.53 (i.e. not bold) to be maximally calibrated. In general, predicting the base rate $\bar{y}$ for all events satisfies the definition of calibration, but lacks boldness. We define Boldness as the spread in probability predictions and measured by
@@ 2.2 Boldness-Recalibration
$(\hat{\delta}_{t},\hat{\gamma}_{t})$ $=argmax_{(\hat{\delta},\hat{\gamma})}(s_{b}:P(M_{c}|\mathbf{y},\hat{\delta},\hat{\gamma})\geq t).$ (9)
Then $\mathbf{x}_{t}$ are called the $(100*t)\%$ boldness-recalibrated predictions where $x_{i,t}=c(x_{i},\hat{\delta}_{t},\hat{\gamma}_{t})$.
The boldness-recalibration methodology formulates finding $\hat{\delta}_{t}$ and $\hat{\gamma}_{t}$ as a nonlinear constrained optimization problem. This optimization is nonlinear because the objective function we optimize is nonlinear with respect to parameters $\delta$ and $\gamma$ and is formulated as
@@
where $x_{i}^{\prime}=c(x_{i},\delta,\gamma)$ and $\bar{x}^{\prime}$ is the mean of $\mathbf{x}^{\prime}$. This is a constrained optimization problem because we restrict the region of parameter values over which we optimize using the following inequality constraint function,
$g(\delta,\gamma)$ $=-(P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})-t)\leq 0,$ (12)
@@
To leverage the efficient gradient-based optimization routines, we also define the Jacobian of (10) as $J_{f}(\delta,\gamma)=\left[\frac{\partial f}{\partial\delta},\frac{\partial f}{\partial\gamma}\right]$ where
$\frac{\partial f}{\partial\gamma}$ $=\frac{-1}{(n-1)sd(\mathbf{x}^{\prime})}\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})\left(log\left(\frac{x_{i}}{1-x_{i}}\right)x_{i}^{\prime}(1-x_{i}^{\prime})-\frac{1}{n}\sum_{j=1}^{n}log\left(\frac{x_{j}}{1-x_{j}}\right)x_{j}^{\prime}(1-x_{j}^{\prime})\right),$ (14)
and the Jacobian of (12) as $J_{g}(\delta,\gamma)=\left[\frac{\partial g}{\partial\delta},\frac{\partial g}{\partial\gamma}\right]$ where
$\frac{\partial g}{\partial\delta}$ $=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\frac{1}{\delta}\left(\sum_{i=1}^{n}\hat{\gamma}_{MLE}^{\prime}(y_{i}-x_{i,MLE}^{\prime})-y_{i}+x_{i}^{\prime}\right),$ (15)
$\frac{\partial g}{\partial\gamma}$ $=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\bigg{(}\sum_{i=1}^{n}log\left(\frac{x_{i}}{1-x_{i}}\right)\big{(}\hat{\gamma}_{MLE}^{\prime}(y_{i}-x_{i,MLE}^{\prime})-y_{i}+x_{i}^{\prime}\big{)}\bigg{)}.$ (16)
Here, $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ are the MLEs for $x_{i}^{\prime}$ and $x_{i,MLE}^{\prime}=c(x_{i}^{\prime};\hat{\delta}_{MLE}^{\prime},\hat{\gamma}_{MLE}^{\prime})$. Full derivation of these gradients can be found in Appendix A.

### v3

@@ Abstract
When probability predictions are too cautious for decision making, boldness-recalibration enables responsible emboldening while maintaining the probability of calibration required by the user. We formulate boldness-recalibration as a nonlinear optimization of boldness with a nonlinear inequality constraint on calibration. We further show that recalibration based on the maximized linear log odds likelihood also maximizes the posterior probability of calibration. We introduce BRcal, an R package implementing boldness-recalibration and supporting methodology as recently proposed. The BRcal package provides direct control of the calibration-boldness tradeoff and visualizes how different calibration levels change individual predictions. We present a new real world case study involving housing foreclosure predictions. The BRcal package is available on the Comprehensive R Archive Network (CRAN) (https://cran.r-project.org/web/packages/BRcal/index.html) and on Github (https://github.com/apguthrie/BRcal).
@@
Acknowledgements The authors are thankful to Leidos for providing funding for this work. The authors would like to thank Matthew Keefe and Andrew McCoy for their insights and roles obtaining case study data for this line of research. The authors are grateful for the helpful comments from the editor and two referees.
@@ 1 Introduction
Boldness-recalibration is a forecaster agnostic approach to responsibly embolden overly cautious probability predictions [1]. When events occur at the rates that were predicted, the probability predictions are well calibrated. Calibration is important so that models provide decision makers with predictions that correspond to that actual rate of events. However, overly cautious predictions tend to satisfy the literal definition of calibration (e.g. predicting sample proportion $\bar{y}$ for all events) but provide little insight on the forthcoming outcome. Predictions that are further from $\bar{y}$ are more actionable. We call predictions that exhibit greater spread (and thus are typically further from $\bar{y}$) more bold and we call the action of spreading predictions out emboldening. To achieve bolder predictions, a sacrifice on calibration is typically required. Boldness-recalibration leverages this tradeoff between calibration and boldness by maximizing boldness at a specified level of calibration, subject to classification accuracy. Rather than rely on calibration assessment techniques such as the Brier Score, Expected Calibration Error (ECE), and Log Loss, we employ a Bayesian model selection approach to quantify the probability of calibration via the posterior model probability. The posterior model probability of calibration is easily interpretable in the context of an individual forecaster, which is an advantage over these common calibration metrics. The key feature of boldness-recalibration that sets it apart from most other calibration techniques is that it provides responsible emboldening of predictions rather than just improving calibration.
In this paper, we present a clear demonstration of how to implement boldness-recalibration in practice and additional mathematical details pertaining to properties of maximum likelihood estimators (MLEs) in [1]. Additionally, we present BRcal: an R package implementing boldness-recalibration and supporting methodology including calibration assessment via Bayesian model selection, LLO-adjustment via MLEs, and visualizations.
@@
The rest of this paper is organized as follows. Section 2 presents a brief and self contained review of the methods presented in [1]. Section 3 presents key mathematical and optimization-related details of maximum likelihood estimation and boldness-recalibration. Section 4 introduces the BRcal package by describing its core functions, implementation, and output from a familiar data example. In Section 5, we demonstrate the capabilities of BRcal by presenting a new example related to housing foreclosure predictions. Section 6 provides concluding remarks.
2 Background
Let $\mathbf{x}$ be a vector of probability predictions from a single source (i.e., a single forecaster, or a single machine learning model, etc) and let $\mathbf{y}$ be a vector of binary event outcomes. We adopt the following likelihood function from [1]
$\pi(\mathbf{y}|\delta,\gamma)=\prod_{i=1}^{n}c(x_{i};\delta,\gamma)^{y_{i}}\left[1-c(x_{i};\delta,\gamma)\right]^{1-y_{i}},$ (1)
@@ 2 Background
and $\delta>0$ and $\gamma\in\mathbb{R}$ are recalibration parameters [23, 24]. Equation (2) maps probability forecasts to the log odds scale, then shifts them by log $\delta$, scales by $\gamma$, and remaps back to the probability scale. Notice the shift is by $log(\delta)$ rather than $\delta$ because the shift occurs on the log odds scale. When we apply (2) to prediction $x_{i}$ with some $\delta$ and $\gamma$, we call this LLO-adjustment.
To assess calibration, we take an approximate Bayesian model selection-based approach comparing a well calibrated model ($\delta$ = $\gamma$ = 1) denoted by $M_{c}$, to an uncalibrated model ($\delta>0$ and $\gamma\in\mathbb{R}$) denoted by $M_{u}$. When $\delta$ = $\gamma$ = 1 is appropriate, no shifting nor scaling of $\mathbf{x}$ is needed, which corresponds to the identify mapping of probability predictions to observed frequencies. This corresponds directly to the frequency-based definition of calibration. The posterior model probability of $M_{c}$ is defined as
$P_{\mathbf{x}}(M_{c}|\mathbf{y})=\frac{1}{1+BF\frac{P(M_{u})}{P(M_{c})}}.$ (3)
where $BF_{\mathbf{x}}$ is the Bayes Factor comparing $M_{u}$ to $M_{c}$ defined as
$BF_{\mathbf{x}}=\frac{P_{\mathbf{x}}(\mathbf{y}|M_{u})}{P_{\mathbf{x}}(\mathbf{y}|M_{c})}.$ (4)
@@
$BF_{\mathbf{x}}\approx exp\left\{-\frac{1}{2}(BIC_{u}-BIC_{c})\right\}.$ (5)
Figure 5 of [1] shows that the BIC approximation for BF performs sensible updates to the posterior odds from the prior odds, even in sample sizes as small as 30. In (5), the BIC under the well calibrated model $M_{c}$ is defined as
@@
$BIC_{u}$ $=2\times log(n)-2\times log(\pi(\hat{\delta}_{MLE},\hat{\gamma}_{MLE}|\mathbf{x},\mathbf{y})),$ (7)
where $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$ are the maximum likelihood estimates (MLEs) for $\delta$ and $\gamma$ in (1). The quantity $P_{\mathbf{x}}(M_{c}|\mathbf{y})$ is the posterior probability $\mathbf{x}$ is calibrated given observed outcomes $\mathbf{y}$ and is used to quantify calibration. MLE recalibration, which is LLO-adjustment of $\mathbf{x}$ using the MLEs, produces maximally calibrated predictions $\mathbf{x}_{MLE}$ where $x_{i,MLE}=c(x_{i};\hat{\delta}_{MLE},\hat{\gamma}_{MLE})$. Section 3.1.2 formally states this result.
Maximally calibrated predictions are not necessarily bold enough for decision making. For example, the Hockey study in [1] showed that even after MLE recalibration, the expert forecaster’s predictions could be more bold while maintaining a high probability of calibration via boldness-recalibration. On the other hand, random noise predictions were reigned in close to the base rate of 0.53 (i.e., not bold) to be maximally calibrated. In general, predicting the base rate $\bar{y}$ for all events satisfies the definition of calibration, but lacks boldness. We define boldness as the spread in probability predictions and measured by
@@
$(\hat{\delta}_{t},\hat{\gamma}_{t})$ $=argmax_{(\delta,\gamma)}(s_{b}:P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y},\delta,\gamma)\geq t),$ (9)
where $x_{i}^{\prime}=c(x_{i},\delta,\gamma)$. Then $\mathbf{x}_{t}$ are called the $(100*t)\%$ boldness-recalibrated predictions where $x_{i,t}=c(x_{i},\hat{\delta}_{t},\hat{\gamma}_{t})$.
3 Methods
3.1 Maximum Likelihood Estimation of $\delta$ and $\gamma$
There is no analytical solution for $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$ as the optimization is mathematically intractable. Thus, all methods in this paper and in [1] rely on numerical optimization routines to acquire solutions for the MLEs. Following common convention, we will minimize the negative log likelihood rather than maximize (1) directly. Additionally, we consider a reparameterization of the negative log likelihood in terms of $\tau=log(\delta)\in\mathbb{R}$. Thus, the MLEs for $\tau$ and $\gamma$ can be written as
$(\hat{\tau}_{MLE},\hat{\gamma}_{MLE})=argmin_{(\tau,\gamma)}\left(-\sum_{i=1}^{n}y_{i}log\left(c(x_{i};exp(\tau),\gamma)\right)+(1-y_{i})log\left(1-c(x_{i};exp(\tau),\gamma)\right)\right),$ (10)
where the MLE for $\delta$ follows directly with $\hat{\delta}_{MLE}=exp(\hat{\tau}_{MLE})$. This reparameterization allows us to optimize in a completely unbounded parameter space, rather than bounding $\delta$ below at 0, which allows greater flexibility in choice of algorithm, as some routines do not allow bound constraints. Additionally, we have found in practice that optimizing in terms of $\tau$ instead of $\delta$ is more efficient, which is explored further in the next section and Appendix E.1.
3.1.1 Optimization Routines for MLEs
In this work, we primarily consider two optimization routines for acquiring MLEs: Nelder-Mead [27] and Limited Memory Broyden–Fletcher–Goldfarb–Shanno (L-BFGS) algorithm [28]. We focus this section on these two algorithms because we have found them to be the most reliable in terms of convergence for MLEs in this work.
When comparing the performance of these two algorithms when used in this work, there is a trade off between efficiency, precision, and stability. In terms of efficiency, L-BFGS has the benefit of being gradient-based. We have found that L-BFGS is 1.06 times faster than Nelder-Mead, on average. However, L-BFGS is 2.00 times faster than Nelder-Mead when closed-form solutions for the gradients of the objective function are provided (on Ubuntu 22.04, R version 4.4.2). The gradients of the objective function being maximized in (10) can be found in Appendix E.1.
The L-BFGS algorithm also benefits from slightly increased precision. When examining the achieved MLEs for the hockey data presented in [1] and in Section 4.1 across a grid of different starting parameter values, the standard deviation in $\hat{\delta}_{MLE}$ from L-BFGS was 3.24 ×$\times 10^{-6}$ for $\hat{\delta}_{MLE}$ and 1.67×$\times 10^{-5}$ for $\hat{\gamma}_{MLE}$. For Nelder-Mead, sd=0.0002 for $\hat{\delta}_{MLE}$ and sd=0.0010 for $\hat{\gamma}_{MLE}$. While Nelder-Mead suffers from a slight deficit in efficiency and precision, we have found it to be more reliable than L-BFGS in that it is less prone to straying from the area in the parameter space where the optimum is located. When L-BFGS iterates away from the optimum, it sometimes produces very extreme parameter values, which can cause the algorithm to abruptly terminate. A more complete comparison of Nelder-Mead, L-BFGS, and a few other readily implemented algorithms can be found in Appendix E.1.
With these tradeoffs in mind, we primarily recommend Nelder-Mead for applications of this work. Nelder-Mead is the default algorithm used in all cases where the MLEs are found in the BRcal package. In cases where speed is of concern, we recommend trying L-BFGS with the closed-form gradients supplied, and in cases where this approach fails, return to Nelder-Mead.
3.1.2 Maximizing $P_{\mathbf{x}^{\prime}}(M_{c}|\mathbf{y})$
In Section 2 and [1], we claim that by LLO-adjusting a set of probability predictions by the MLEs for $\delta$ and $\gamma$, their calibration will be maximized. This result can be more formally written as follows.
Result 1.
Let $\mathbf{x}$ be a set of probability predictions with corresponding outcomes $\mathbf{y}$. Let $\mathbf{x}^{\prime}$ be $\mathbf{x}$ under LLO-adjustment by some $\delta$ and $\gamma$. Then under the BIC approximation of $BF_{\mathbf{x^{\prime}}}$,
$(\hat{\delta}_{MLE},\hat{\gamma}_{MLE})=argmax_{(\delta,\gamma)}P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y}),$
meaning MLE recalibration maximizes the posterior probability of calibration.
In other words, the MLEs, which by definition maximize the LLO likelihood, also maximize the posterior probability of calibration as considered here. See Appendix D for proof of this result.
3.2 Boldness-Recalibration
The problem of solving (9) to get the boldness-recalibration parameter estimates is also mathematically intractable. In this paper, we formulate finding $\hat{\delta}_{t}$ and $\hat{\gamma}_{t}$ as a nonlinear constrained optimization problem.
Again following common optimization conventions, we will minimize some objective function and enforce that some constraint function be less than or equal to zero. Since the goal of boldness-recalibration is to maximize boldness, measured by $s_{b}$, we will minimize negative $s_{b}$. Thus, our objective function is as follows
@@ 3.2 Boldness-Recalibration
where $\bar{x}^{\prime}$ is the mean of $\mathbf{x}^{\prime}$. Since boldness-recalibration requires that $(P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})\geq t)$ we have the following constraint function after rearranging to follow the convention mentioned above
$g(\delta,\gamma)$ $=-(P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})-t)\leq 0,$ (13)
where $P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})$ is the probability $\mathbf{x}^{\prime}$ is calibrated given $\mathbf{y}$. For clarity in describing and plotting the constraint, we refer to $P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})$ as the constraint surface, as it has the same shape and optimum as (13). Notice both the objective function $f(\delta,\gamma)$ and constraint function $g(\delta,\gamma)$ are nonlinear with respect to parameters $\delta$ and $\gamma$.
@@
3.2.1 Nonlinear Optimization Routines for Boldness-Recalibration
Since the constraint in (13) is nonlinear with respect to $\delta$ and $\gamma$, we must consider different optimization routines than discussed in Section 3.1.1. In this work, we will consider routines that are readily available in NLopt [29], a free open source nonlinear optimization library.
For all sets of predicted probabilities we encountered or generated for this work, the constraint surface (the posterior model probability) has had one unique maximum located at the MLEs (shown in Appendix D). Datasets include real-world case studies and simulated data intentionally generated with various levels of miscalibration. The constraint region tends to take on a single connected region for fixed levels of $t$. Additionally, the objective function (the negative standard deviation) tends to have one unique minimum within the constraint region for fixed levels of $t$. Neither of these results are mathematically guaranteed, however, we have found them to be true in all cases we have explored, even in the most extreme cases of miscalibration. A single connected region with a unique optimum is helpful for convergence, as it ensures that the optimizer will not get stuck in a disjoint, subpar region of the parameter space that satisfies the constraint.
NLopt includes both global and local optimization routines. Since we are typically operating within a single constraint region with a single optimum, we need not consider global routines that explore the entire space. To ensure the optimizer can locate the region of interest, we recommend starting the optimization at the MLEs, since they are at the center of the region, as they maximize $P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})$.
NLopt also includes both gradient-based and derivative-free routines. Given that the partial first derivatives of (11) and (13) are available in closed-form, we will consider gradient-based optimization routines, as they are typically more efficient. We define the Jacobians of (11) and (13) in Appendices A.3 and A.4, respectively.
With this in mind, we chose to focus on local gradient based approaches of which there were three options that supported non-linear inequality constraints: Sequential Least-Squares Quadratic Programming (SLSQP) [30, 31], Augmented Lagrangian algorithm (AUGLAG) [32, 33], and Method of Moving Asymptotes (MMA) [34]. Of these, we found AUGLAG produced the most stable results in terms of convergence in boldness-recalibration problems and is used throughout this work.
The implementation of AUGLAG provided by nloptr reformulates the constrained optimization problem as a two-stage optimization. The inner stage optimizes the objective function with a penalty for violating the constraint rather than use the constraint directly. The outer stage governs the penalty. If the inner stage returns a value that violates the constraint despite the penalty, the outer stage imposes a larger penalty. Note that since the inner stage is itself an unconstrained optimization routine, users may specify any optimization routine that supports nonlinear objective functions. For the purpose of this package, we found that SLSQP as the inner optimization routine in AUGLAG provided the most stable results, thus SLSQP is set as the default inner optimization routine. For a more complete comparison of optimization routines and subroutines, see Appendix E.2.
The original proposal of boldness-recalibration in [1] does not explain how to practically implement the technique. Early implementations of boldness-recalibration used a grid search-based approach to find $\hat{\delta}_{t}$ and $\hat{\gamma}_{t}$. However, this approach is inefficient and lacks precision. Compared to the original naive grid search approach, the nonlinear constrained optimization approach described above is 139 times faster, on average (on Ubuntu 22.04, R version 4.4.2).